OpenStudy (anonymous):

Which point lies in the solution set for the following system of inequalities? y > x + 4 y > -2x + 2 A. (-4, 0) B. (-3, 2) C. (2, -1) D. (0, 6)

jimthompson5910 (jim_thompson5910):

Let's check choice A y > x + 4 0 > -4 + 4 ... plug in x = -4 and y = 0 0 > 0 ... which is FALSE so choice A is NOT the answer

jimthompson5910 (jim_thompson5910):

you do the same for each choice check to see if the point satisfies both inequalities if it makes both inequalities true, then it's the answer

jimthompson5910 (jim_thompson5910):

tell me what answer you get so I can check it

OpenStudy (anonymous):

I got C, is that correct?

jimthompson5910 (jim_thompson5910):

no C is not correct, here's why y > x + 4 -1 > 2 + 4 ... plug in x = 2, y = -1 -1 > 6 that is FALSE as well, so C is NOT the answer

OpenStudy (anonymous):

I tried all of them and none worked..?

jimthompson5910 (jim_thompson5910):

y > x + 4 6 > 0 + 4 ... plug in (0,6) 6 > 4 .... which is TRUE ---------------------- y > -2x + 2 6 > -2(0) + 2 ... plug in (0,6) 6 > 0 + 2 6 > 2 ... which is also TRUE

jimthompson5910 (jim_thompson5910):

do you see how I'm getting this?

OpenStudy (anonymous):

oh, yes I get it now I was putting the numbers in backwards o_o thank you!

jimthompson5910 (jim_thompson5910):

oh that might make things a bit more complicated lol, glad you figured out what went wrong